25. Analyzing a Portfolio [LO2, 4] You have $100,000 to invest in a portfolio containing Stock X and Stock Y. Your goal is to with formula please
Answer #1

Expected return of the portfolio is computed using the below formula:

Rp= Wx*Rx + Wy*Ry

0.1270= Wx*0.114 + (1 - Wx)*0.0868

0.1270= 0.114Wx + 0.0868 - 0.0868Wx

0.1270 - 0.0868= 0.00272Wx

0.00272Wx= 0.0402

Wx= 0.0402/ 0.0272

= 1.4779*100

= 147.79%.

Wy= 1 - 1.4779

= -0.4779*100

= -47.79%

Money invested in stock Y= 0.4779*$100,000

= $47,790.

Beta of the portfolio:

= 1.4779*1.25 + (-0.4779*0.85)

= 1.8474 - 0.4062

= 1.4412.

In case of any query, kindly comment on the solution.